OpenStudy (anonymous):

Brass is made from a mixture of copper and other elements. A mixture that is 80% copper is combined with a mixture that is 60% copper, resulting in 100 pounds of brass that is 65% copper. Which equation can be used to find x, the amount of 60% mixture used to create the 65% mixture? a. 0.8(100 – x) + 0.6x = 100(0.65) b. 0.6(100 – x) + 0.8x = 100(0.65) c. 0.8(100) + 0.6x = 0.65(100 - x) d. 0.6(x) + 0.8(100 + x) = 0.65

OpenStudy (anonymous):

@mathgeek27

OpenStudy (anonymous):

give me a sec :)

OpenStudy (anonymous):

I say C, but you should get a second opinion bc I'm not 100% sure @thadyoung what's your opinion?

OpenStudy (anonymous):

@cody7744 Is C correct ?

OpenStudy (anonymous):

looks right

OpenStudy (shinalcantara):

let 'x' be the amount @ 60% 'y' be the amount @ 80% 0.80y + 0.60x = 100(0.65) ----equation 1 since x and y combined is 100 then x + y = 100 y = 100 - x ---equation 2 substitute equation 2 to 1 0.80(100 - x) + 0.60x = 100(0.65)

OpenStudy (anonymous):

So A ?

OpenStudy (shinalcantara):

yep